# Fuss-Free Network: A Simplified and Efficient Neural Network for Crowd Counting

### v1

@@ 3.3 Focus Transition Module
where F is the input, C(·) is the channel feature weight optimization, Y(·) is the dynamic convolution, P(·) represents the output after stacked dynamic convolutions, S(·) is the spatial feature weight optimization, $\bar{P}$ is the mean value of P, and MLP is the multi-layer perceptron.
The focus transition module we designed will fully unleash the potential of the streamlined framework, aiding the model in extracting richer, more targeted semantic information, achieving smooth transition effects, and enhancing the accuracy and robustness of crowd counting tasks.
3.4 Loss Function
@@
The total loss consists of three components: count loss, optimal transport(OT) loss [20], and variation loss.
Count loss is a critical loss in crowd counting that quantifies the gap between the number of people predicted by the model and the actual number of people labeled, thus contributing to better convergence of the model. It is defined as the formula 6. Based on the counting loss, we further constructed a variation loss using the predicted density map and the ground truth density map. It employs the L1 norm to measure the intensity or grayscale variation differences between adjacent pixels in the predicted and ground truth density maps. Through this approach, our goal is to enhance the smoothness of the image and effectively reduce the adverse effects of potential noise on density estimation. Its definition is as follows:
$\mathcal{L}_{c}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|c_{p}-c_{g}\right|,\mathcal{L}_{v}=\frac{1}{N}\sum_{i=1}^{N}\left|x_{i}-y_{i}\right|,$ (6)
where $c_{p}$ is the predicted population and $c_{g}$ is the actual population, N represents the batch size, $x_{i}$ represents the density value of the $i$-th pixel in the true density map, and $y_{i}$ represents the predicted density value of the $i$-th pixel at the corresponding position in the predicted density map.
@@
OTLoss is used to measure the difference between two probability distributions, making it particularly suitable for handling misaligned data distribution problems. It encourages the model to learn probability densities that are closer to the true distribution. It first calculates the distance matrix between the predicted density map and the true density map using the Euclidean distance. Then, it utilizes the Sinkhorn algorithm for iterative optimization, finding the optimal transportation plan from the source distribution to the target distribution, and obtaining the dual variables during this process. Finally, based on these dual variables and the predicted density distribution, OTLoss is computed. This loss measures the discrepancy between the predicted density distribution and the true labels in terms of optimal transportation. Throughout the entire process, the regularization parameter regulates the balance between distance cost and distribution smoothness. Its complete definition is as follows:
$D_{ij}$ $=\sqrt{\left(x_{i}-x_{j}\right)^{2}+\left(y_{i}-y_{j}\right)^{2}},$ (7)
where each matrix element $D_{ij}$ represents the Euclidean distance between the $i$-th grid point in the predicted density map, with coordinates ($x_{i}$, $y_{i}$), and the $j$-th ground-truth point in the true density map, with coordinates ($x_{i}$, $y_{j}$).
$\mathbf{\beta}=\log(\mathbf{P})-\frac{1}{\lambda}\left[\log\left(\sum_{j=1}^{N}e^{-\lambda D_{ij}}\right)\right]_{i=1}^{M},$ (8)
where $\beta$ represents the dual variables obtained through the Sinkhorn algorithm, $D$ is the distance matrix, $P$ represents the prediction map, $M$ represents the number of rows in matrix $D$, NN represents the number of columns, and $D_{ij}$ represents the value at the $i$-th row and $j$-th column of the distance matrix, and $\lambda$ is the regularization parameter set to 10.
$\mathcal{L}_{ot}$ $=\sum_{i=1}^{N}\left[\left(\frac{\mathbf{\beta}_{i}}{P_{i}}-\frac{\mathbf{\beta}_{i}\cdot\mathbf{\rho}_{i}}{P_{i}^{2}}\right)\mathbf{\rho}_{i}\right],$ (9)
where $N$ represents the batch size, $\beta_{i}$ is the dual variable computed by the Sinkhorn algorithm for the $i$-th sample, $P_{i}$ is the sum of predicted densities in the $i$-th sample, and $\rho_{i}$ is the vector composed of all predicted density values for the $i$-th sample.
Finally, the total loss is the weighted sum of three components. Let $\lambda_{1}$ and $\lambda_{2}$ be the weighting values, the total loss is defined as:
$\mathcal{L}_{all}$ $=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{1}+\lambda_{2}\mathcal{L}_{2},$ (10)
Combining counting loss, OTLoss, and variation loss can enhance the accuracy and robustness of crowd counting models. Additionally, regularization reduces noise in the predicted maps, improves model generalization, accelerates convergence, and helps train more accurate and stable crowd counting models. In our experiments, we fixed $\lambda_{1}$ at 0.1 and only fine-tuned $\lambda_{2}$.
4 Experiment
This section validates the effectiveness of our method on crowd counting benchmarks through extensive experiments and ablation studies, comparing it with state-of-the-art techniques.
4.1 Implementation Details
4.1.1 Dataset
@@
We evaluate our methods in four benchmarks, including UCF_CC_50 [29], ShanghaiTech Part A and B [6], and NWPU-Crowd [30]. These datasets differ in terms of scene complexity, crowd density range, and image resolution.
UCF_CC_50 [29] consists of 50 images of high-density crowds, with a total of 63,075 individuals across the entire dataset. The number of individuals in each image ranges from 94 to 4543, with an average of 1280 individuals per image. Extreme variations in crowd density and occlusion situations are among the challenges presented by the densely crowded complex scenes in this dataset. Due to the relatively small size of the dataset, Idrees et al. [29] defined a cross-validation protocol for training and validating models.
ShanghaiTech [6] comprises a total of 1198 annotated images with 330,165 individually marked heads, divided into Part A and Part B. Part A consists of 482 images, among which 300 are allocated for training and 182 for testing; these images were sourced from the Internet. In contrast, Part B comprises 400 training images and 316 testing images, all sourced from real surveillance footage. These two parts of the dataset, due to their complex real-life scenarios, diverse environmental conditions, and distribution of crowds, pose significant challenges for crowd counting technology. They are crucial benchmarks for measuring and improving the accuracy of crowd counting algorithms.
NWPU-Crowd [30] differs from the previous ones in several aspects. It has a larger scale, containing 5,109 high-resolution images covering various lighting conditions, perspectives, and scene types. The number of people ranges from tens to tens of thousands, providing more diverse scenes and a wider range of people counts. Additionally, the dataset includes a few negative samples that contribute to the model’s increased robustness. Compared to the previous datasets, it imposes higher precision requirements and tests the generalization ability of crowd counting algorithms.
4.1.2 Training Settings and Hyperparameters
@@
For a fair comparison, the backbone of FFNet is the official ConvNeXt-Tiny model, which is pretrained on the ImageNet-1k dataset. In our experiments, data augmentation was limited to random cropping and random horizontal flipping. For ShanghaiTech Part B, we set the cropping size to 512; for NWPU, the cropping size is set to 384; and for UCF_CC_50 and ShanghaiTech Part A datasets, we adjust the cropping size to 256. The model training utilized the AdamW optimizer with a batch size of 8, and the initial learning rate was set to 1e-5. Additionally, to mitigate overfitting, we introduced an L2 regularization term with a coefficient of 0.005. All experiments were conducted using PyTorch on a single 16-GB RTX 4080.
4.1.3 Evaluation Metrics
In crowd counting research, Mean Absolute Error (MAE) and Mean Squared Error (MSE) are commonly used as evaluation metrics for crowd counting. The Mean Absolute Error (MAE) is the average of the absolute differences between the predicted counts and the actual counts. It provides an intuitive measure of how much the predicted values deviate from the actual values. The Mean Squared Error (MSE) is the average of the squared differences between the predicted counts and the actual counts. It is highly sensitive to outliers and requires the model to have good robustness to noise and outliers in the data. The expressions for MAE and MSE are as follows:
$MAE$ $=\frac{1}{N}\sum_{i=1}^{N}\left|c_{i}-\hat{c}_{i}\right|,$ (11)
$MSE$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(c_{i}-\hat{c}_{i}\right)^{2}},$ (12)
@@ 4.1.3 Evaluation Metrics
where N is the number of test images, and $c_{i}$ and $\hat{c}_{i}$ are the predicted and true number of people in the ith image, respectively.
@@
Method Venue Backbone Parameters FLOPS UCF_CC_50 SHA SHB NWPU(V) MAE MSE MAE MSE MAE MSE MAE MSE MCNN [6] CVPR16 CNN - - 377.6 509.1 110.2 173.2 26.4 41.3 218.5 700.6 CSRNet [2] CVPR18 VGG-16 - - 266.1 397.5 68.2 115.0 10.6 16.0 104.9 433.5 CAN [10] CVPR19 VGG-16 - - 212.2 243.7 62.3 100.0 7.8 12.2 93.5 489.9 SFCN [31] CVPR19 ResNet-101 - - 214.2 318.2 64.8 107.5 7.6 13.0 - - BL [20] ICCV19 VGG-19 - - 229.3 308.2 62.8 101.8 7.7 12.7 93.6 470.4 ASNet [15] CVPR20 VGG-16 - - 174.8 251.6 57.8 90.1 - - - - AMRNet [32] ECCV20 VGG-16 - - 184.0 265.8 61.5 98.3 7.0 11.0 - - DM-Count [21] NeurIPS20 VGG-19 - - 211.0 291.5 59.7 95.7 7.4 11.8 70.5 357.6 P2PNet [23] ICCV21 VGG-16 19.22M 104.71G 172.7 256.2 52.7 85.1 6.3 9.9 - - DKPNet [33] ICCV21 HRNet-W40 12.60M 46.31G - - 55.6 91.0 6.6 10.9 61.8 438.7 SASNet [38] AAAI21 VGG-16 38.90M 232.98G 161.4 234.46 53.5 88.3 6.3 9.9 - - MAN [16] CVPR22 VGG-19 30.96M 105.85G - - 56.8 90.3 - - - - CHFL [34] CVPR22 VGG-19 21.51M 108.27G - - 57.5 94.3 6.9 11.0 76.8 343.0 GauNet [35] CVPR22 ResNet-50 - - 186.3 256.5 54.8 89.1 6.2 9.9 - - CLTR [36] ECCV22 ResNet-50 40.95M 28.73G - - 56.9 95.2 6.5 10.6 61.9 246.3 FIDT [40] TMM22 HRNet-W48 66.58M 142.40G 171.4 233.1 57.0 103.4 6.9 11.8 51.4 107.6 PML [22] TCSVT23 ConvNeXt-S - - - - 53.4 93.1 6.2 9.7 - - LDFNet [37] arXiv23 VGG-19 - - 164.9 234.8 52.7 82.9 6.7 11.5 - - PET [19] ICCV23 VGG-16 51.65M 859.68G - - 49.3 78.9 6.2 9.7 58.5 238.0 STEERER [13] ICCV23 HRNet-W48 64.42M 94.40G - - 54.5 86.9 5.8 8.5 54.3 238.3 BCCT [17] arXiv21 T2T-ViT-14 - - - - 53.1 82.2 7.3 11.3 53.0 170.3 CCTrans [8] arXiv21 Twins-SVT 103.58M 99.44B 168.7 234.5 52.3 84.9 6.2 9.9 38.6 87.8 CUT [18] BMVC22 Twins-PCPVT 68.09M 90.45G - - 51.9 79.1 - - - - FFNet(ours) - ConvNeXt-T 29.02M 23.67G 161.1 233.3 48.3 80.5 6.1 9.0 41.2 113.2
@@
Table 1: Compared with the advanced methods on UCF_CC_50, ShanghaiTech Part A and B, and NWPU datasets. And compare the parameter quantity and flops with advanced methods. The best results are shown in bold, and the second-best results are underlined.
4.2 Comparison with state-of-the-art methods
4.2.1 Counting performance
@@
CNN-based crowd counting methods dominate the field of crowd counting, with most approaches building upon the foundation of CNNs for model design. As shown in Table 1, our method outperforms them, exhibiting significant advantages in several benchmark tests and achieving state-of-the-art (SOTA) or near-state-of-the-art performance levels. Even when compared to the popular transformer-based methods in recent years, our model demonstrates superiority, achieving better results in most cases.
@@
Specifically, we achieved outstanding results on the UCF_CC_50 dataset, with an MAE of 161.4, surpassing existing methods and reaching a new state-of-the-art level. On the ShanghaiTech Part A subset, we demonstrated significant superiority with an MAE value of 48.3, successfully surpassing the latest SOTA in this field. In the evaluation of the ShanghaiTech Part B subset, despite fierce competition, our method still achieved a highly competitive second-place performance with an MAE of 6.1. Finally, in the more challenging NWPU-Crowd dataset, our method also achieved results close to the top tier. Without excessive complexity in design, our model achieves incredible results on these datasets. This fully demonstrates the effectiveness of FFNet.
4.2.2 Parameters and Computational Resources
@@
Adhering to the original intention of our work, we aimed to simplify the model while achieving high crowd counting performance. Therefore, we conducted lightweight testing on FFNet and state-of-the-art methods. The testing primarily focused on open-source models released in the past three years. We did not test some methods that are partially open-source or not open-source. For a fair comparison, we all test with an all-one tensor of size (1, 3, 512, 512) as input. And we measured the lightweight characteristics of the model based on two metrics: model parameters and floating-point operations (FLOPs). Among them, the model parameters is one of the key indicators for measuring the complexity of neural network models. It refers to the total number of weights and bias parameters in the model. Models with fewer parameters occupy less storage resources, have faster inference speeds, are easier to deploy, and have lower risks of overfitting. FLOPS is another important metric for measuring the complexity and computational requirements of neural network models. In the experiment, it refers to the number of floating-point operations required for one forward pass. Models with lower FLOPS consume fewer computational resources, have faster inference speeds, are more suitable for low-power environments, and are easier to deploy on mobile and embedded devices.
@@
The experimental testing results are presented in Table 1. The data shows that over the past three years, as counting performance has continued to improve, there has been a significant increase in model parameters and FLOPs, resulting in greater model complexity. However, FFNet’s parameters and FLOPs are significantly lower than those of other advanced crowd counting methods. While achieving superior counting performance, achieve optimal FLOPs and highly competitive model parameters. We believe that FFNet’s SOTA counting performance with such low FLOPs and model parameters is truly commendable. This effectively proves the efficacy of our proposed model: concise, efficient, low in complexity, and yet powerful in performance. FFNet effectively balances model simplification and performance enhancement, making it suitable not only for deployment in resource-constrained environments but also for performing exceptionally well in scenarios with high real-time requirements. This significantly drives the development and application of efficient crowd-counting technologies.
4.3 The Effectiveness of Our Simplified Structure
To validate the effectiveness of our concise network framework, we conducted corresponding experiments. We replaced FTM with ordinary convolutions. Specifically, we replaced FTM with four concatenated simple convolutions on top of a simple framework.
Method SHA SHB
MAE MSE MAE MSE
SASNet [38] 53.5 88.3 6.3 9.9
CCTrans [8] 52.3 84.9 6.2 9.9
CUT [18] 51.9 79.1 - -
PET [19] 49.3 78.9 6.2 9.7
STEERER [13] 54.5 86.9 5.8 8.5
FFNet+CNN 51.7 89.0 6.2 9.5
FFNet+FTM 48.3 80.5 6.1 9.0
Table 2: Simple network experiment: A comparison between a simple network with ordinary convolutions and a simple network with FTM against state-of-the-art methods.
The results, as shown in Table 2, demonstrate that, compared to state-of-the-art approaches, our framework achieved an MAE of 51.7 on the ShanghaiTech Part A dataset and 6.2 on the Part B dataset, solely utilizing ordinary convolutions for intermediate downsampling and feature extraction. This surpasses most state-of-the-art methods. This further confirms the effectiveness of the streamlined network architecture we adopted and verifies the positive role of the FTM module in enhancing the overall network performance.
4.4 Multi-scale fusion experiment

### v3

@@ II-C Focus Transition Module
where $\mathbf{F}$ is the input, $C(\cdot)$ is the channel feature weight optimization, $Y(\cdot)$ is the dynamic convolution, $P(\cdot)$ represents the output after stacked dynamic convolutions, $S(\cdot)$ is the spatial feature weight optimization, $\mathbf{\bar{P}}$ is the mean value of $\mathbf{P}$, and MLP is the multi-layer perceptron.
The fundamental concept behind dynamic convolution is to dynamically combine multiple parallel convolution kernels based on their attentions, which are determined by the input data. This approach generates convolution kernel weights during the runtime of the model, allowing for adaptability to different input data. By capturing features across various dimensions such as input channels, output channels, kernel size, and the number of kernels, dynamic convolution offers flexibility in feature extraction. Dynamic convolution utilizes a combination of average pooling and convolution to determine attention weights, which are then dynamically adjusted. This allows for adaptive parameter adjustment of the convolutional kernel to better align with the input data’s features. We define $\mathbf{X}$ as the input of dynamic convolution and the process of dynamic convolution attention extraction and dynamic weighting as follows:
where $A(\cdot)$ represents the attention calculation, AvgPool denotes average pooling, FC stands for fully connected layer, Conv indicates convolution operation, n represents the number of convolutional kernels, and $\sigma$ is the sigmoid activation function; $Y(\cdot)$ represents the dynamic convolution weighting process, $\mathbf{A}_{nk}$, $\mathbf{A}_{sk}$, $\mathbf{A}_{ik}$, $\mathbf{A}_{ok}$ are the attention weights on four dimensions: the number of convolution kernels, kernel size, input channels, and output channels, respectively, $\mathbf{W}_{k}$ represents the weights of different convolution kernels, and C denotes the input optimized by weights.
II-D Loss Function
@@
We employed the Loss function utilized in DM-Count [14] which consists of three components: count loss, optimal transport (OT) loss [25], and variation loss.
Count loss is a critical loss in crowd counting that quantifies the gap between the number of people predicted by the model and the actual number of people labeled, thus contributing to better convergence of the model. It is defined as follows:
$\mathcal{L}_{c}(\mathbf{z},\hat{\mathbf{z}})$ $=\left|\|{{\mathbf{z}}}\|_{1}-\|{\hat{\mathbf{z}}}\|_{1}\right|,$ (6)
where $\mathbf{z}$ represents the vectorized binary map for dot-annotation and $\hat{\mathbf{z}}$ is the vectorized predicted density map returned by network, $\|\cdot\|_{1}$ denotes the L1 norm.
@@
OTLoss is used to measure the difference between two probability distributions, making it particularly suitable for handling misaligned data distribution problems. It encourages the model to learn probability densities that are closer to the true distribution. It utilizes the Sinkhorn algorithm [26] for iterative optimization, finding the optimal transportation plan from the source distribution to the target distribution, and obtaining the dual variables during this process. Finally, based on these dual variables and the predicted density distribution, OTLoss is computed. Its definition is as follows:
$\mathcal{L}_{ot}(\hat{\mathbf{z}})$ $=\langle\frac{\mathbf{\beta}^{\ast}}{\|{\hat{\mathbf{z}}}\|_{1}}-\frac{\langle\mathbf{\beta}^{\ast},\hat{\mathbf{z}}\rangle}{\|{\hat{\mathbf{z}}}\|^{2}_{1}},\hat{\mathbf{z}}\rangle,$ (7)
[…]
Furthermore, variation loss is utilized to measure the intensity or grayscale variation differences between adjacent pixels in the predicted and ground truth density maps. Specifically, the L1 norm is employed. We utilize this loss function to improve the overall image smoothness and effectively minimize the negative impact of potential noise on density estimation. To balance the contribution of losses from areas with different crowd densities, we use the actual crowd count as a weighting factor to adjust the variance loss. Its definition is as follows:
$\mathcal{L}_{v}(\mathbf{z},\hat{\mathbf{z}})$ $=\|{{\mathbf{z}}}\|_{1}\cdot\||\mathbf{z}-\hat{\mathbf{z}}\|_{1}.$ (8)
The total loss is the weighted sum of three components, which is defined as:
$\mathcal{L}_{all}$ $=\mathcal{L}_{c}+\lambda_{1}\mathcal{L}_{ot}+\lambda_{2}\mathcal{L}_{v},$ (9)
where $\lambda_{1}$ and $\lambda_{2}$ are the weighting values, and were assigned as 0.1 and 0.01, respectively, in the experiment.
III Experiments
III-A Implementation Details
III-A1 Dataset
@@
We evaluate our methods in four benchmarks, including UCF_CC_50 [27], ShanghaiTech Part A (SHA) and ShanghaiTech Part B (SHB) [4], and NWPU-Crowd [28]. The datasets, described in detail in TableI, vary in terms of scene complexity, crowd density range, and image resolution.
Datasets Numbers of Image Average Resolution Total marked persons Max marked persons Min marked persons Average marked persons UCF_CC_50 50 2101 × 2888 63974 4633 94 1279 ShanghaiTech A(SHA) 482 589 × 868 241677 3139 33 501 ShanghaiTech B(SHB) 716 768 × 1024 88488 578 9 123 NWPU 5109 2191 × 3209 2133375 20033 0 418
TABLE I: Crowd counting datasets used for evaluation.
III-A2 Training Settings and Hyperparameters
@@
For a fair comparison, the backbone of FFNet is the official ConvNeXt-Tiny model, which is pretrained on the ImageNet-1k dataset. In our experiments, data augmentation was limited to random cropping and random horizontal flipping. For SHB, we set the cropping size to 512; for NWPU, the cropping size is set to 384; and for UCF_CC_50 and SHA datasets, we adjust the cropping size to 256. The model training utilized the AdamW optimizer with a batch size of 8, and the initial learning rate was set to 1e-5. Additionally, to mitigate overfitting, we introduced an L2 regularization term with a coefficient of 0.005. All experiments were conducted using PyTorch on a single 16-GB RTX 4080.
III-A3 Evaluation Metrics
In crowd counting research, Mean Absolute Error (MAE) and Mean Squared Error (MSE) are commonly used as evaluation metrics for crowd counting. The MAE and MSE are defined as follows:
$MAE$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{c}_{i}-\hat{\mathbf{c}}_{i}\right|,$ (10)
$MSE$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\mathbf{c}_{i}-\hat{\mathbf{c}}_{i}\right)^{2}},$ (11)
@@ III-A3 Evaluation Metrics
where $N$ is the number of test images, and $\mathbf{c}_{i}$ and $\hat{\mathbf{c}}_{i}$ are the true and predicted number of people in the $i$-th image, respectively.
@@
Method Venue Backbone Parameters FLOPS UCF_CC_50 SHA SHB NWPU(V) MAE MSE MAE MSE MAE MSE MAE MSE CAN [7] CVPR19 VGG-16 - - 212.2 243.7 62.3 100.0 7.8 12.2 93.5 489.9 ASNet [11] CVPR20 VGG-16 - - 174.8 251.6 57.8 90.1 - - - - DensityCNN [29] TMM20 VGG-16 - - 244.6 341.8 63.1 106.3 9.1 16.3 - - DM-Count [14] NeurIPS20 VGG-19 - - 211.0 291.5 59.7 95.7 7.4 11.8 70.5 357.6 P2PNet [16] ICCV21 VGG-16 19.22M 104.71G 172.7 256.2 52.7 85.1 6.3 9.9 - - DKPNet [30] ICCV21 HRNet-W40 12.60M 46.31G - - 55.6 91.0 6.6 10.9 61.8 438.7 SASNet [19] AAAI21 VGG-16 38.90M 232.98G 161.4 234.46 53.5 88.3 6.3 9.9 - - PFDNet [31] TMM21 VGG-16 - - 205.8 289.3 53.8 89.2 6.5 10.7 - - MAN [12] CVPR22 VGG-19 30.96M 105.85G - - 56.8 90.3 - - - - CHFL [32] CVPR22 VGG-19 21.51M 108.27G - - 57.5 94.3 6.9 11.0 76.8 343.0 CLTR [33] ECCV22 ResNet-50 40.95M 28.73G - - 56.9 95.2 6.5 10.6 61.9 246.3 FIDT [34] TMM22 HRNet-W48 66.58M 142.40G 171.4 233.1 57.0 103.4 6.9 11.8 51.4 107.6 PML [15] TCSVT23 ConvNeXt-S - - - - 53.4 93.1 6.2 9.7 - - PET [17] ICCV23 VGG-16 51.65M 859.68G - - 49.3 78.9 6.2 9.7 58.5 238.0 STEERER [9] ICCV23 HRNet-W48 64.42M 94.40G - - 54.5 86.9 5.8 8.5 54.3 238.3 BCCT [5] arXiv21 T2T-ViT-14 - - - - 53.1 82.2 7.3 11.3 53.0 170.3 CCTrans [18] arXiv21 Twins-SVT 103.58M 99.44G 168.7 234.5 52.3 84.9 6.2 9.9 38.6 87.8 CUT [6] BMVC22 Twins-PCPVT 68.09M 90.45G - - 51.9 79.1 - - - - FFNet(ours) - ConvNeXt-T 29.02M 23.67G 161.1 233.3 48.3 80.5 6.1 9.0 41.2 113.2
@@
TABLE II: Compared with the advanced methods on four datasets. (The best results are shown in bold, and the second-best results are underlined).
III-B Comparison with state-of-the-art methods
III-B1 Counting performance
@@
CNN-based crowd counting methods dominate the field of crowd counting, with most approaches building upon the foundation of CNNs for model design. As shown in Table II, our method outperforms them, exhibiting significant advantages in several benchmark tests and achieving SOTA or near-state-of-the-art performance levels. Even when compared to the popular transformer-based methods in recent years, our model demonstrates superiority, achieving better results in most cases.
@@
Specifically, we achieved outstanding results on the UCF_CC_50 dataset, with an MAE of 161.4, surpassing existing methods and reaching a new SOTA level. On the SHA subset, we demonstrated significant superiority with an MAE value of 48.3, successfully surpassing the latest SOTA in this field. In the evaluation of the SHB subset, despite fierce competition, our method still achieved a highly competitive second-place performance with an MAE of 6.1. Finally, in the more challenging NWPU-Crowd dataset, our method also achieved results close to the top tier. Without excessive complexity in design, our model achieves excellent results on these datasets.
III-B2 Parameters and Computational Resources
@@
Adhering to the original intention of our work, we aimed to simplify the model while achieving high crowd counting performance. Therefore, we conducted lightweight testing on FFNet and SOTA methods. The testing primarily focused on open-source models released in the past three years. We did not test some methods that are partially open-source or not open-source. For a fair comparison, we all test with an all-one tensor of size (1, 3, 512, 512) as input. And we measured the lightweight characteristics of the model based on two metrics: model parameters and floating-point operations (FLOPs). Among them, the model parameters is one of the key indicators for measuring the complexity of neural network models. It refers to the total number of weights and bias parameters in the model. Models with fewer parameters occupy less storage resources, have faster inference speeds, are easier to deploy, and have lower risks of overfitting. FLOPS is another important metric for measuring the complexity and computational requirements of neural network models. In the experiment, it refers to the number of floating-point operations required for one forward pass. Models with lower FLOPS consume fewer computational resources, have faster inference speeds, are more suitable for low-power environments, and are easier to deploy on mobile and embedded devices.
@@
The experimental testing results are presented in Table II. The data shows that over the past three years, as counting performance has continued to improve, there has been a significant increase in model parameters and FLOPs, resulting in greater model complexity. However, FFNet’s parameters and FLOPs are significantly lower than those of other advanced crowd counting methods. While achieving superior counting performance, achieve optimal FLOPs and highly competitive model parameters. This effectively proves the efficacy of our proposed model: concise, efficient, low in complexity, and yet powerful in performance. FFNet effectively balances model simplification and performance enhancement, making it suitable not only for deployment in resource-constrained environments but also for performing exceptionally well in scenarios with high real-time requirements.
III-C The Effectiveness of Our Simplified Structure
To verify the effectiveness of our concise structure, we conducted replacement experiments on the backbone of the structure. We replaced the backbone of our model, CCTrans, and PET, with different pre-trained networks. Subsequently, we performed parameter fine-tuning on the modified models mentioned above using the SHA dataset, and conducted a count of the number of parameters and Flops for each model after replacing the backbone, using a 512×512 input. The results are shown in Table III. When using ConvNeXt-Tiny and Swin-Transformer-Small, our structure achieved significantly lower MAE and MSE with a minimal number of parameters and floating-point operations compared to the other CCTrans and PET. In addition, when using MobileNet-v3 as the backbone, our structure achieved MAE and MSE that are second only to CCTrans, with the lowest number of parameters (3.05M) and floating-point operations (1.51G). The experimental results show that our structure exhibits better robustness and effectiveness with lower parameter and computational complexity compared to other the structures of CCTrans and PET.
Method MobileNet-v3 ConvNeXt-Tiny Swin-Transformer-Small Parameters FLOPS MAE MSE Parameters FLOPS MAE MSE Parameters FLOPS MAE MSE CCTrans [18] 4.74M 9.08G 62.9 102.1 32.12M 41.01G 53.9 86.5 37.35M 47.78G 56.6 92.4 PET [17] 15.19M 133.81G 66.6 110.6 90.96M 343.54G 54.1 85.1 106.68M 404.46G 58.1 90.0 FFNet 3.05M 1.51G 64.3 109.3 29.02M 23.67G 48.3 80.5 34.26M 30.44G 54.7 88.1
TABLE III: Backbone replacement experiment. (Comparision of the effectiveness of different structs.)
We also compare our model with recent lightweight crowd counting models using MobileNet-v3 as the backbone. As show in Table IV, by simply using a lightweight backbone, our model easily achieves lightweight performance.
Method Parameters FLOPS(input size) SHA
MAE MSE
ECCNAS-Lat [35] 3.79M 7.38G(512×512) 66.3 116.5
ECCNAS-Acc [35] 3.88M 16.46G(512×512) 62.0 110.9
LMSFFNet [36] 4.58M 14.9G(1024×768) 85.8 139.96
FFNet 3.05M 1.51G(512×512) 64.3 109.3
(MobileNet-v3) 4.53G(1024×768)
TABLE IV: Comparison with advanced lightweight models. (We utilized 256×256 cropped images to calculate the MAE and MSE metrics for the models.)
III-D Multi-scale feature fusion